I'm looking for the name of a theorem, and the completion of the theorem. I've been searching the internet and my algebra books, but can't locate the theorem. The question marks is what I need to fill in. If you know the theorem, I can look it up to complete the statement.
The ? Theorem states if p and q are real numbers and if p^2 = q^2, then ?
The ? Theorem states if p and q are real numbers and if p * q = 0, then ?
Thanks!
Answer by Edwin McCravy(20081) 
 I'm looking for the name of a theorem, and the completion of the theorem. I've been searching the internet and my algebra books, but can't locate the theorem. The question marks is what I need to fill in. If you know the theorem, I can look it up to complete the statement.
The ? Theorem states if p and q are real numbers and if p^2 = q^2, then ?
The ? Theorem states if p and q are real numbers and if p * q = 0, then ?
Thanks!
I can answer only three of the four:
The ????? Theorem states if p and q are real numbers and if p^2 = q^2, then
|p| = |q|
The zero-factor Theorem states if p and q are real numbers and if p * q = 0,
then either p=0 or q=0 or both.
Sorry, but I do not know of any name for the first. My guess would be
the "absolute value theorem", but that's only a guess.
Edwin
|
|
|
| |